OpenStudy (anonymous):

Please help! How would I solve 12 5/8 + 9 2/3?

OpenStudy (anonymous):

Is that 12 and 5/8 or (125)/8, also 9 and 2/3 or (92)/3?

OpenStudy (anonymous):

12 and 5/8, with 9 and 2/3

OpenStudy (tux):

\[\frac{ 8*12+5 }{ 8 }+\frac{ 3*9+2 }{ 3 }\]=Your turn

OpenStudy (anonymous):

do you just add the denominators though? Like they dont need to be the same?

OpenStudy (tux):

Multiply the whole number part by the fraction's denominator. Add that to the numerator Write that result on top of the denominator.

OpenStudy (tux):

Here you can see the steps http://www.wolframalpha.com/input/?i=101%2F8%2B29%2F3